Summary of Evening shift 9/4/99

Patricia McBride (mcbride@fnpsph)
Sun, 05 Sep 1999 00:13:37 CDT

Shift Summary for Eve shift, 09.04.99

On Shift: P. McBride, Sasha Glasov

Also present: Elliott, Jason L

Hours of beam delivered at all intensities: 8.5
Hours of beam delivered at nominal intensity: 8.5
Hours of E832 data (special runs not included): 8.25
Hours of special runs (laser scans, muon runs...): 0.2

Accounting of lost time: kfrends
Runs ended this shift:

14428 BEAM_832 22 Million events
14429 Muon

Tapes brought to FCC this shift: none

Accesses during shift: None

Overview of shift:

Stable beam for most of the shift; many small problems

Description of Shift:

run spill time comment
- - - ------------------------------------------------------------------------
14428 15:35 page Ed to ask him how to restart DAQ monitoring
process. He says it is complicated and we
can wait until the end of the run.

Called MCR to ask about beam conditions; Kevin
says everything is fine; will check BOA.

15:45 Elliott appears and shows us how to
restart DAQ monitoring process with the script

BobT arrives with a fresh supply of coffee

26-30 no triggers; reset pipeline and did a kfrend

106 onsplit request for tape QKA024, mounted tape

17:30 second onsplit request, mounted tape QKCC18

18:00 paged JasonH to check on extra frowny face on
etot#sum plot. He said ``Don't worry about it''

18:15 DaveN stops by to say that everything is fine
with the gas system.

140 18:20 ktevana errors --> kfrend

18:20 third onsplit request, mounted QKN034

143-144 18:23 more ktevana errors --> 2 kfrends (trigger in same
error state
after first one)

Page Ed to ask about error message daqdead_sec = 1-2
which are popping up about every fifth spill.
Error often seems to be correlated with long
processing time on KTeV4. Paged Jason Ladue to ask
to remove the extra bananas from the readout at the
end of this run in order to reduce the overall event

185-191 19:20 Trigger is in an error state; reload kumquats and
bananas. Do a few more kfrends.

196 19:34 ktevana errors - 2 kfrends

213-214 19:55 ktevana errors - 2 more kfrends

218-219 no beam

220 the beam is back, but SEM = 6.99e12 and remains high
(6.9 - 7.0e12) for about 15 spills

231 20:11 ktevana errors - 1 kfrend

236 20:26 ktevana errors - 1 kfrend

278 21:25 ktevana errors - 1 kfrend

21:30 fourth onsplit request; problems mmounting KQNL32

308 22:05 ktevana errors - 1 kfrend
KQNL32 mounts successfully after Sasha cleans the

309 22:06 ktevana errors - 1 more kfrend

22:10 Paged JasonL to tell him that tapes are 90% full
He is on his way...

317 22:18 SEM is reading close to 7e12 again
Called MCR to ask them to reduce it to 6.8e12

22:29 ktevana errors - 1 kfrend; JasonL arrives

335 22:41 Ed calls, at the same moment a voice says that the
tapes are full:
end run; start to setup for muon run,
change tapes, call MCR; push crash button; Jason
bananas for 1x and and 2x from the readout to reduce
event size

22:50 fifth onsplit request; it is for QKSM17 this time

22:51 MCR calls to say that magnets are ready for muon mode

14429 23:00 Muon run started

23:04 ftt write problems with QKSM17 drive tps3d2 on KTeV2;
Sasha cleans this drive too; Jim keeps on complaining;
Cleaning helps, QKSM17 mounts successfully

5 23:06 Stop muon run (250k events); setup for hadrons
switch magnet NM4AN to positive
Jason removes remaining X view fruit from the readout

14430 23:18 BEAM_832 run started with a reset of bananas and
and a kfrend; lost the permit in stream 5; kfrend

6 23:21 running smoothly; livetime is at 0.75; no daq_dead
errors for
about 20 spills

24-26 23:50 low intensity for a few spills (2e12)

24:00 3.993 Million Vacuum 2pi0

Follow up Notes on Previous Shift

Items to Follow up on Next Shift